### v1

# Hier-SLAM++: Neuro-Symbolic Semantic SLAM with a Hierarchically Categorical Gaussian Splatting

Abstract
We propose Hier-SLAM++, a comprehensive Neuro-Symbolic semantic 3D Gaussian Splatting SLAM method with both RGB-D and monocular input featuring an advanced hierarchical categorical representation, which enables accurate pose estimation as well as global 3D semantic mapping. The parameter usage in semantic SLAM systems increases significantly with the growing complexity of the environment, making scene understanding particularly challenging and costly. To address this problem, we introduce a novel and general hierarchical representation that encodes both semantic and geometric information in a compact form into 3D Gaussian Splatting, leveraging the capabilities of large language models (LLMs) as well as the 3D generative model. By utilizing the proposed hierarchical tree structure, semantic information is symbolically represented and learned in an end-to-end manner. We further introduce a novel semantic loss designed to optimize hierarchical semantic information through both inter-level and cross-level optimization. Additionally, we propose an improved SLAM system to support both RGB-D and monocular inputs using a feed-forward model. To the best of our knowledge, this is the first semantic monocular Gaussian Splatting SLAM system, significantly reducing sensor requirements for 3D semantic understanding and broadening the applicability of semantic Gaussian SLAM system. We conduct experiments on both synthetic and real-world datasets, demonstrating superior or on-par performance with state-of-the-art NeRF-based and Gaussian-based SLAM systems, while significantly reducing storage and training time requirements.
Index Terms:
Semantic SLAM, hierarchical category, gaussian splatting, RGB-D, monocular.
I Introduction
Visual Simultaneous Localization and Mapping (SLAM) is a key technology for ego-motion estimation and scene perception, widely employed in applications across various domains, such as autonomous drones [1], self-driving vehicles [2], and interactive experiences in Augmented Reality (AR) and Virtual Reality (VR) [3]. Semantic information, which provides high-level knowledge about an environment, is indispensable for comprehensive scene understanding. Thus, it plays a crucial role in enabling intelligent robots to perform various complex tasks. In recent decades, advances in image segmentation and enhanced map representations have accelerated progress in semantic visual SLAM [4, 5, 6, 7, 8, 9, 10, 11, 7, 12].
Figure 1: (a). The hierarchical tree is generated by integrating geometric information and semantic messages, utilizing 3D generative models and Large Language Models (LLMs). (b). The global 3D Gaussian map generated by Hier-SLAM++ with learned semantic labels is shown on the left. The established hierarchical tree of the semantic information is organized on the right. Based on the established tree, the hierarchical symbolic representation for semantic information is shown at the bottom of the block, which compresses semantic data, reducing both memory usage and training time of the semantic SLAM. (c). The rendered semantic map at different levels shows a coarse-to-fine understanding, beneficial for real-world scenarios with shifting perspectives from distant to close.
Recently, 3D Gaussian Splatting has gained attention as a leading 3D world representation method [13, 14, 15], thanks to its fast rendering and optimization efficiency enabled by the highly parallelized rasterization of 3D primitives. Specifically, 3D Gaussian Splatting models the continuous distributions of geometric parameters through Gaussian distributions. This approach not only improves the quality of 3D representation but also supports efficient optimization, making it particularly well-suited for SLAM tasks. While promising, current SLAM systems based on 3D Gaussian Splatting [16, 17, 18, 19, 20] primarily focus on geometric reconstruction. However, the lack of semantic integration limits its potential for complex tasks such as visual navigation, planning, and autonomous driving. Addressing this gap by incorporating semantic understanding without sacrificing efficiency remains a key challenge.
A naive solution would be to augment 3D primitives with discrete semantic labels and parameterize their distribution using a categorical distribution, i.e., a flat Softmax-based representation. However, 3D Gaussian Splatting is already a storage-intensive approach [21, 22], requiring a large number of 3D primitives with multiple parameters to achieve realistic rendering. Adding semantic distribution parameters would further increase storage requirements and processing time, with grow linear in the number of semantic classes, making it particularly impractical for complex scene understanding. Furthermore, given a fixed number of observations (video frames), increasing the number of optimizable parameters leads to the curse of dimensionality, resulting in degraded optimization performance in high-dimensional spaces, such as when handling a large number of semantic classes. Overall, this added overhead for semantic learning is especially problematic for robotics applications, where real-time performance and limited resources are critical. Recent studies have explored alternative methods to mitigate these challenges. For instance, [23] directly learns a 3-channel RGB visualization to avoid semantic labels learning. [24] adopts a flat semantic representation and relies on a pre-trained foundation model to generate an embedding feature map rather than an explicit 3D semantic label map.
Unlike flat representations, semantic information naturally organizes into a hierarchical structure of classes, as shown in Fig. 1. The attributes (such as structures) and properties (such as surfaces) of the semantic classes are extracted as symbolic nodes to generate a hierarchical tree, where each semantic class is represented as a hierarchical expression based on a root-to-leaf path. This hierarchy enable efficient coding of extensive information with a relatively small number of symbolic nodes to learn, resulting in a compact representation. For example, a binary tree with a depth of $10$ can represent $2^{10}$ classes, allowing for the encoding of 1,024 classes using only 20 codes (i.e., $2\times 10$ through 2-dimensional Softmax coding at each level), enabling a reduction of up to $O(\log N)$. Building upon this concept, Hier-SLAM [25] was first introduced as an RGB-D Gaussian Splatting SLAM that employs a hierarchical tree to represent semantic information in an unknown environment, achieving strong performance with improved storage efficiency and operational speed.
Hier-SLAM [25] have demonstrated the advantages and potential of incorporating hierarchical categorization for a semantic RGB-D SLAM system. However, it is limited by a simplistic hierarchical tree representation and lacks robust semantic optimization. To address these limitations and further improve its effectiveness, we propose Hier-SLAM++, a comprehensive neuro-symbolic semantic Gaussian Splatting SLAM framework that incorporates an advanced hierarchical representation for semantic information. Different from Hier-SLAM, we incorporate both geometric (i.e., shape and size details) and semantic information into the hierarchical symbolic tree expression using both Large Language Models (LLMs) and 3D Generative models. Specifically, we utilize a text-to-3D generative model [26] to extract shape information for each class and organize it into hierarchical levels. Additionally, we leverage LLMs to include both semantic and size information in the tree structure, effectively addressing the lack of size information in 3D generative models. With the established tree, semantic information in the environment is represented as symbolic nodes along root-to-leaf paths, to be learned in an end-to-end manner during the SLAM operation. To enhance semantic understanding, we introduce an improved hierarchical loss function that incorporates both inter-level and cross-level optimizations. This coarse-to-fine hierarchical approach based on the established hierarchical representation is beneficial for real-world applications, especially those involving observations from distant to nearby views. Furthermore, our SLAM system is extended to support both RGB-D and monocular inputs. We utilize the geometric output of the 3D feed-forward model [27] as a geometric prior, removing the depth dependency of existing SLAM methods. This enables Hier-SLAM++ to perform semantic SLAM without relying on dedicated depth sensors, broadening the range of applicable scenarios. Finally, we enhance and refine the 3D Gaussian Splatting SLAM to improve both performance and efficiency. We conduct experiments on multiple synthetic and real-world datasets. The results demonstrate that our SLAM system outperforms or is on par with state-of-the-art NeRF-based and Gaussian-based SLAM methods. Additionally, our approach effectively reduces storage and training time.
The technical contributions of our work include:
1) We introduce an advanced hierarchical tree representation that integrates semantic, shape, and size information, leveraging the capabilities of both LLMs and 3D generative models. This tree-based coding efficiently compresses semantic information while maintaining its inherent hierarchical structure, resulting in reduced memory usage and shorter training times.
2) We present a novel optimization approach for the proposed semantic hierarchical representation, integrating both inter-level and cross-level losses to ensure thorough optimization across all levels of the hierarchical semantic information.
3) We propose an improved semantic SLAM system capable of supporting both RGB-D and monocular inputs by leveraging a 3D feed-forward model. To the best of our knowledge, Hier-SLAM++ is the first Gaussian Splatting based semantic SLAM system to support monocular input, significantly reducing sensor requirements and broadening its application scope.
4) We conduct extensive experiments on both synthetic and real-world datasets, comprehensively demonstrating the effectiveness of our method and its improvements.
II Related Work
II-A Visual SLAM system
Visual Simultaneous Localization and Mapping (SLAM) using visual sensors is a long-standing and active research topic that focuses on the simultaneous estimation of ego-motion and the reconstruction of a global 3D map. Traditional SLAM methods [28, 29, 30, 31], developed over many years, have demonstrated both high accuracy and robustness, even in challenging environments. This makes them an indispensable frontend foundation for various real-world applications, such as autonomous driving and drones. In recent years, several new 3D representations have emerged in the 3D computer vision fields. With the advent of neural implicit representations, numerous works [32, 33, 34, 35] have demonstrated the effectiveness of NeRF-based SLAM approaches. These new representations introduce novel capabilities to the SLAM community, including enhanced rendering capability. 3D Gaussian Splatting is the latest advancement in the 3D vision field, offering significantly faster rendering speed and strong performance compared with neural implicit representations. Recent works [16, 17, 20, 19] have demonstrated new SLAM capabilities by integrating 3D Gaussian Splatting techniques, further pushing the boundaries of real-time scene understanding and reconstruction.
II-B Neural implicit visual SLAM
Neural Radiance Fields (NeRF) [36] is a neural rendering technique that represents 3D scenes using a fully connected neural network. iMap [32] proposes a real-time keyframe-based RGB-D SLAM method that uses a multilayer perceptron (MLP) as the sole scene representation. Nice-SLAM [33] introduces a dense SLAM system that incorporates multi-level local information by employing a hierarchical scene representation. Nicer-SLAM [34] extends this approach by proposing a monocular SLAM system with a hierarchical neural implicit map representation, eliminating the dependency on depth information. Integrating SLAM with semantic understanding, many works [7, 11, 12] have explored neural implicit representations for semantic mapping and localization tasks. DNS-SLAM [7] incorporates 2D semantic priors with a coarse-to-fine geometry representation to embed semantic information into the map. SNI-SLAM [11] integrates geometry, appearance, and semantic features into a shared feature space to enhance the robustness of the SLAM method. However, these approaches are constrained by the slow convergence of neural implicit map representations, which often leads to performance degradation when coupled with semantic objectives [37, 38]. In contrast, Gaussian Splatting provides significant advantages, including fast rendering performance and high-density reconstruction quality, making it a promising alternative for map representation.
II-C 3D Gaussian Splatting SLAM
3D Gaussian Splatting has recently gained attention as a promising 3D representation due to its fast rendering performance and high-density reconstruction quality. Leveraging this 3D representation, SplaTAM [16] uses silhouette guidance for pose estimation and map reconstruction in RGB-D SLAM systems, while MonoGS [17] supports both monocular and RGB-D SLAM. Photo-SLAM [19] utilizes 3D points generated by ORB-SLAM as position parameters for Gaussian Splatting primitives within the SLAM system. These studies highlight the strong potential of 3D Gaussian Splatting in SLAM tasks across various scenarios [16, 17, 18, 19, 20]. However, integrating semantic understanding into SLAM presents a significant challenge, as it requires jointly solving three high-dimensional optimization problems with varying value ranges and convergence behaviors, while also significantly increasing storage requirements. Focus on the integration of semantic information into the Gaussian Splatting SLAM, SGS-SLAM [23] added RGB 3-channels to create a semantic visualization map to avoid true semantic labels learning. SemGauss-SLAM [24] uses a flat semantic representation supervised by a large pre-trained foundation model. However, these methods overlook the natural hierarchical structure of the real world. Additionally, relying on large foundation models increases the complexity of the neural network and its computational demands, with performance largely dependent on the embeddings from these pre-trained models. Eliminating the dependency on foundation models, this work proposes a simple yet effective hierarchical representation for semantic understanding in the SLAM system which largely reduces the training storage and time.
III Method
Figure 2: Left: Overview of the Hier-SLAM++ pipeline. The global 3D Gaussian map is initialized using the first frame of the video stream input. The system then alternates between the Tracking and Mapping steps as new frames are processed (see Section III-D). In the RGB-D setting, depth is directly obtained from sensor input, whereas in the monocular setting the 3D feed-forward method (DUSt3R) is used to generate a geometric prior for depth estimation (see Section III-C). Top Right: Hierarchical representation of semantic information. The Tree Generation process leverages the capabilities of both LLMs and 3D Generative Models in a top-to-bottom manner. This hierarchical tree is used to establish a symbolic coding for each Gaussian primitive (see Section III-A). Additionally, we introduce a novel loss function that combines Inter-level Loss $L_{\text{Inter}}$ and Cross-level Loss $L_{\text{Cross}}$ to optimize the hierarchical semantic representation (see Section III-B). Bottom Right: An example of hierarchical semantic rendering.
The entire pipeline of our method is illustrated in Fig. 2. Within Hier-SLAM++, we model the semantic space as a hierarchical tree, where semantic information is represented as a root-to-leaf hierarchical symbolic structure. This representation not only comprehensively captures the hierarchical attributes of semantic information, but also effectively compresses storage requirements, while enhancing semantic understanding with valuable scaling-up capability. To construct an effective tree, we leverage both LLMs and 3D Generative Models to integrate semantic and geometric information. Next, each 3D Gaussian primitive is augmented with a hierarchical semantic embedding code, which can be learned in an end-to-end manner through the proposed Inter-level and Cross-level semantic losses. Furthermore, our method is extended to support monocular input by incorporating geometric priors extracted from the feed-forward method, thereby eliminating the dependency on depth information. The following sections provide detailed descriptions for each part of our Hier-SLAM++.
III-A Hierarchical representation
We propose hierarchical symbolic representation for semantic information to provide improved efficiency, compression, and compactness for semantic SLAM, making it well-suited for multi-parameter joint optimization and enabling scalability to complex scenes. In this section, we provide a detailed explanation of the hierarchical tree parametrization and its generation process.
III-A1 Tree Parametrization
The hierarchical tree $\bm{G}$ adopted in Hier-SLAM++ is represented as the set of vertices and edges, $\bm{G}=(\bm{V},\bm{E})$. The vertices set $\bm{V}=\cup_{l=0}^{L}\left\{\bm{v}_{l}\right\}$ comes from the classes from all tree levels, where $\left\{\bm{v}_{l}\right\}$ represents the set of nodes at the $l$-th level of the tree, and we use $L$ to represent depth of the whole tree. The edge set $\bm{E}=\cup_{m=0}^{L-1}\left\{\bm{e}_{m}\right\}$ captures the subordinate relationships, reflecting semantic attribution, size message, and geometric knowledge. For the $i$-th semantic class $g^{i}$, which is treated as a single leaf node in the tree view, its hierarchical expression is:
$g^{i}=\{v_{l}^{i},e_{m}^{i}\mid l=0,1,...,L;\;m=0,1,...,L-1\},$ (1)
which corresponds to the root-to-leaf path: $g^{i}=v_{0}^{i}\xrightarrow{e_{0}}v_{1}^{i}\xrightarrow{e_{1}}\cdots\xrightarrow{e_{L-2}}v_{L-1}^{i}\xrightarrow{e_{L-1}}v_{L}^{i}$, as illustrated in Fig. 3. Take a semantic class $g=\texttt{Bed}$ as an example, a 5-level tree symbolic coding can be represented as: $\left\{v_{0}:\texttt{LargeItems}\right\}\rightarrow\left\{v_{1}:\texttt{Furnishings}\right\}\rightarrow\left\{v_{2}:\texttt{BedroomFurniture}\right\}\rightarrow\left\{v_{3}:\texttt{Rectangular}\right\}\rightarrow\left\{v_{4}:\texttt{Bed}\right\}$. The hierarchical attributes and properties of class $g^{i}$ are represented by the symbolic node information $v_{l}^{i}$, while its relationships such as ‘include’ and ‘possessing’ are represented by the edge information $e_{m}^{i}:\rightarrow$. In this way, every semantic class can be coded in a progressive, symbolic hierarchical manner, incorporating both semantic and geometric perspectives. Moreover, the standard flat representation can be seen as a single-level tree coding from the hierarchical viewpoint. Herein, we shall omit the superscripts (class indexes) of variables for clarity and compact notation.
III-A2 Tree Generation with LLMs and 3D Generative Models
To represent semantic information both accurately and comprehensively, an effective tree structure should incorporate information from both semantic and geometric perspectives, as these are the two crucial types of information in an unknown environment, which are essential for achieving a comprehensive understanding of the global 3D environment. For capturing semantic meanings, such as functional relationships and summary description labels, we use Large Language Models (LLMs), specifically GPT-4 Turbo [39], known for its strong commonsense reasoning and language understanding capabilities. On the other hand, to capture general geometric information without dataset limitations, we employ 3D Generative Models, specifically text-to-3D models, as they have the capability to generate generalized and unified 3D objects with accurate shape information based on text prompts after training. We adopt the text-to-3D approach MeshGPT [26], which learns 3D symbolic embeddings with the local mesh geometry and topology of 3D assets, making it well-suited for our purpose. However, the text-to-3D method only generates 3D objects within a unified coordinate system, omitting size information. To address this issue, we complement the geometric shape information with the size attributes derived from LLMs, to capture the comprehensive geometric information for each semantic class. The leverage of both LLMs and 3D Generative Models enables the creation of an elaborate and well-structured hierarchical tree representation, where each level reflects a distinct aspect of a class’s attributes, enhancing the depth and utility of the global 3D understanding. The following part provides a detailed explanation of the tree generation process.
As illustrated in Fig. 3, the tree generation process follows a top-to-bottom manner. At the upper level, the initial grouping divides all classes into several subgroups based on their typical physical sizes in scenes, obtained through prompt queries to the LLM:
$\{v,e\}_{0}=\mathcal{L}^{\text{size}}\{g\}$ (2)
where $\mathcal{L}^{\text{size}}$ represents the LLMs usage for the size message, and the subscript in $\{v,e\}_{0}$ represents the generation at the initial levels. The clustering results categorize classes into groups such as small items, medium items, and large items. We provide details of the specific prompt used for the LLM in Appendix 1.
Next, the classes within each group are used as input prompts to LLMs for further clustering based on their semantic relationships, particularly their functional properties:
$\{v,e\}_{1}=\mathcal{L}^{\text{func}}\{g\}$ (3)
where $\mathcal{L}^{\text{func}}$ denotes the LLMs clustering based on functional relationships, and the subscript of $\{v,e\}_{1}$ stands for the following levels generation. We demonstrate the detail prompt used for the LLM in Appendix 1.
After multiple levels generation using LLMs, 3D Generative Models are employed to cluster classes based on their geometric shapes. Specifically, we adopt MeshGPT [26] to capture the geometric information for each semantic class. MeshGPT [26] is a text-to-3D model that encodes object shape and internal topology relationships into latent quantized embeddings, forming a symbolic 3D representation in the shape space. Given a text prompt (i.e., the class name) as input to the pre-trained MeshGPT, we extract the latent quantized embeddings to represent the shape information for the class:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{z}=\mathcal{E}\{g\}\in\mathbb{N}^{F\times 2}}}$ (4)
where $g$ is the class name provided as input. $\mathcal{E}$ represents the pre-trained MeshGPT encoder, and $\bm{z}$ is the quantized geometric embedding extracted from the pre-trained encoder. For each object’s embedding, $F$ is the number of triangle faces used in the mesh representation. Since the number of triangle faces varies across different semantic objects, i.e., different values of $F$, we first compute the average shape embedding $\bm{z}_{\text{avg}}\in\mathbb{R}^{1\times 2}$ for each object. The averaged embeddings from all classes are then concatenated and subsequently normalized to a standard distribution (zero mean and unit variance) to ensure consistency across classes:
$\tilde{\bm{z}}=\mathcal{N}\{\mathcal{C}(\bm{z}_{\text{avg}})\}$ (5)
where $\mathcal{N}$ and $\mathcal{C}$ denote the normalization and concatenation operations, respectively. Next, the normalized embeddings $\tilde{\bm{z}}$ are clustered using K-means++, denoted by the operator $\mathcal{K}$. To address the lack of descriptive labels for each clustering result, we further employ LLMs to summarize and refine the clustering outcomes:
$\{v,e\}_{2}=\mathcal{L}^{\text{sum}}\{\mathcal{K}(\tilde{\bm{z}})\}$ (6)
where $\mathcal{L}^{\text{sum}}$ represents the use of LLMs for summarization. The generated descriptive labels, such as ”box”, ”flat”, ”soft”, summarize the common geometric attributes within each grouping result, further enhancing the interpretability of the hierarchical tree. The details are illustrated in Appendix 1.
Whenever LLMs are employed, we ensure balanced, meaningful, and descriptive outputs through carefully designed prompt inputs, as detailed in Appendix 1. Furthermore, an additional LLM validator [25] is utilized after each generation level to verify clustering completeness and accuracy, ensuring that no classes are missing or incorrectly included (see Appendix 1 for detail). Finally, a global tree evaluation is conducted to detect and resolve any duplicates or omissions across tree levels, ensuring a fully balanced and complete hierarchical tree structure. We use manual inspection as the final verification to ensure the correctness and rationality of the holistically generated tree.
Figure 3: Visualization of the hierarchical tree structure and semantic embedding. For tree representation, each semantic class is expressed hierarchically as $g=\{v_{l},e_{m}\}$, where edges $e_{m}$ are depicted as lines connecting tree nodes. The nodes within the same level are illustrated via the same color. For tree generation, we utilize both LLMs and 3D Generative Models to extract and group messages. For hierarchical semantic embedding, we represent semantic information using a root-to-leaf symbolic path. We propose two types of representations: the one-hot representation and the binary representation, which can reduce the original dimensionality by up to $O(\log^{2}N)$.
III-B Hierarchical optimization
Based on the generated hierarchical tree structure, the semantic information is represented as compact hierarchical symbolic nodes, and further integrated into the 3D Gaussian primitives, learned in an end-to-end manner with the input frame stream. To ensure accurate semantic understanding, we introduce a semantic hierarchical loss, which consists of Inter-level and Cross-level losses. These losses are incorporated into the optimization process during the online SLAM operation. The following sections provide a detailed explanation of the proposed tree encoding and the loss calculation.
III-B1 Tree Encoding
As introduced in Eq. (1), the semantic class in Hier-SLAM++ is represented hierarchically in a root-to-leaf symbolic manner based on the generated tree. This effectively compacts semantic information, facilitates SLAM joint optimization, and provides valuable scalability for semantic understanding in unknown environments. Specifically, we propose two types of hierarchical semantic embedding approaches, both of which compact the original semantic information.
One-hot representation: As demonstrated in Fig. 3, the hierarchical semantic embedding $\bm{h}$ is composed of the one-hot embeddings across all levels:
$\bm{h}=\mathcal{C}(\bm{h}_{l})\in\mathbb{B}^{N},\quad\bm{h}_{l}\in\mathbb{B}^{n}$ (7)
where $l$ represents the $l$-th level of the tree, and $\mathcal{C}$ denotes the concatenation operation. For each level’s representation $\bm{h}_{l}$, we use $n$-dimensional one-hot codes (in colored boxes) to encode the symbolic nodes (in colored circles) from the tree of this level, as shown in Fig. 3. The dimension $n$ corresponds to the maximum number of nodes at the $l$-th level. Using the one-hot representation, the overall dimension of the semantic embedding is the sum of the dimensions across all levels, given by $N=\sum n$. This enable a maximum reduction of up to $O(\log N)$ compared to the original dimension $N_{\text{original}}$ in the flat representation. This hierarchical representation effectively compacts the semantic embedding, reducing storage requirements and facilitating joint optimization.
Binary representation: We propose a further compact version of our hierarchical semantic representation in this section. The binary representation of the hierarchical semantic embedding $\bm{b}$ consists of the binary embeddings $\bm{b}_{l}$ across all tree levels:
$\bm{b}=\mathcal{C}(\bm{b}_{l})\in\mathbb{B}^{K},\quad\bm{b}_{l}\in\mathbb{B}^{k}$ (8)
where $\bm{b}_{l}$ is the binary representation converted from the one-hot semantic embedding $\bm{h}_{l}$ at each level: $\bm{h}_{l}\rightarrow\bm{b}_{l}$, as depicted in the last row example in Fig. 3. For instance, if the one-hot embedding has a dimension of $n=8$, this corresponds to a maximum of $8$ unique codes at this level (covering values from 00 to $7$). After binarization, the dimension of the compact version is reduced to $k=3$, since $82^{3}=$ 8. By leveraging binary representation, the hierarchical coding dimension at each level is reduced to $k=\log n$, compared to the one-hot encoding with dimension $n$. Consequently, the total semantic encoding dimension across all levels is given by $K=\sum k$, resulting in a more compact representation.
III-B2 Loss Calculation
To fully optimize the hierarchical semantic message both within each level and across different levels of the tree, the proposed hierarchical loss consists of Inter-level loss $L_{\text{Inter}}$ and Cross-level loss $L_{\text{Cross}}$:
$L_{\text{{Semantic}}}=\omega_{1}L_{\text{{Inter}}}+\omega_{2}L_{\text{{Cross}}}$ (9)
where $\omega_{1}$ and $\omega_{2}$ are balancing coefficients for each loss term. Specifically, the Inter-level loss $L_{\text{Inter}}$ is employed within each level’s embedding $\bm{h}^{l}$ to ensure the optimization of semantic information within each level:
$L_{\text{{Inter}}}=\sum_{l=0}^{L}L_{\text{{ce}}}(\mathcal{S}(\bm{h}^{l}),\mathcal{P}^{l})$ (10)
where $L_{\text{{ce}}}$ indicates the cross-entropy semantic loss, and $\mathcal{P}^{l}$ denotes the semantic ground truth at level $l$. $\mathcal{S}$ stands for the Softmax operation, converting embeddings into probabilities. For our proposed compact version with binary representation $\bm{b}^{l}$, we replace the original cross-entropy loss with binary cross-entropy loss $L_{\text{bce}}$ for the binary Inter-level loss calculation:
$L_{\text{Inter}}^{\text{bin}}=\sum_{l=0}^{L}L_{\text{{bce}}}(\mathcal{S}(\bm{b}^{l}),\mathcal{P}^{l})$ (11)
To enforce accurate understanding across tree levels, we introduce the Cross-level loss, which is computed using a semantic decoder based on the holistic hierarchical codes $\bm{h}$. Specifically, we first perform a 2D convolution operation on the semantic embedding, following an activation layer (i.e., ReLU). Then, a second 2D convolution layer is applied to map the hidden embedding into the flat coding, followed by a softmax operation to convert the embeddings into probabilities. The overall Cross-level loss $L_{\text{Cross}}$ is defined as:
$L_{\text{Cross}}=L_{\text{ce}}\big{(}\mathcal{S}(\text{decoder}(\bm{h})),\mathcal{P}\big{)},$ (12)
where $\mathcal{P}$ represents the semantic ground truth, and the definition of $\text{decoder}(\bm{h})$ is:
$\text{decoder}(\bm{h})=\text{Conv}\big{(}\text{ReLU}(\text{Conv}(\bm{h}))\big{)}$ (13)
III-C Monocular setting with geometric priority
Accurate depth information provides valuable geometric cues for global map reconstruction and pose estimation in a SLAM system. In an RGB-D setting, depth signals are directly obtained from sensors. However, in a monocular setting without depth information, SLAM [28, 29, 30, 31] typically relies on triangulation to initialize the global mapping. Yet, due to the inherent coupling between pose estimation and mapping, this approach often suffers from relatively lower accuracy compared with RGB-D based methods. In our Hier-SLAM++, we employ a feed-forward method as a geometric prior. Specifically, we leverage DUSt3R [27] to provide depth geometric priors for monocular SLAM. DUSt3R is able to generate both a 3D point map and pose estimations from several sparse-view image inputs, which can serve as the geometric prior for the monocular SLAM. The following section provides details on the monocular setting in our Hier-SLAM++.
III-C1 Geometric prior acquisition
DUSt3R [27] supports sparse view inputs with a minimum of two views. Larger view inputs enhance reconstruction accuracy by providing more information, yet at the cost of increased computational demands. To balance reconstruction accuracy and computational efficiency, we sample every $S$ frames with a skip of $\delta$ frames from the input video to form an image set $\{I_{i}\}$.
$\{I_{i}\;|\;i=i_{0}+\delta\cdot(s-1),\;s=1,\dots,S\}$ (14)
Each image set is then fed into DUSt3R to generate geometric information for every image frame, specifically the depth maps. Although DUSt3R is also able to provide pose predictions for each input frame, we observe that the accuracy of estimated pose is relatively low—performing even worse than a constant velocity pose prediction in the SLAM system based on our experiments. Therefore, only the estimated geometric information from DUSt3R is leveraged in our method.
III-C2 Geometric prior correction
The scale of the estimation for each image set varies and also differs from the scale of the online global map estimation during SLAM operation, requiring an additional geometric prior correction operator.
Firstly, we calculate the initial scale and shift parameters by comparing the rendering results from the current global map prediction with the depth prior obtained from DUSt3R. Specifically, the depth map $D_{c}$ and the silhouette map $S_{c}$ for current frame $I_{c}$ are rendered based on the global Gaussian map $G$ as well as the pose estimation $\mathbf{T}_{c}$ for the current frame:
$\{D_{c},S_{c}\}=\mathcal{R}(G,\mathbf{T}_{c})$ (15)
where $\mathcal{R}$ represents the rasterization procedure for the 3D Gaussian map. Next, the global scale and shift parameters $\lambda_{c},\tau_{c}$ are calculated by comparing the predicted depth $\hat{D}_{c}$ from DUSt3R with the rendered depth $D_{c}$ using the least squares estimation on the visible regions $M_{c}$ defined by $S_{c}$:
$\{\lambda_{c},\tau_{c}\}=\arg\min\left(\hat{D}_{c}^{\text{align}}-D_{c}\right)^{2}M_{c}$ (16)
where $\hat{D}_{c}^{\text{align}}=\lambda_{c}\cdot\hat{D}_{c}+\tau_{c}$.
Secondly, the estimated parameters are fed into the optimization as initialization and further optimized jointly within the map optimization via the depth supervision loss:
$\{\lambda_{c},\tau_{c},G\}=\arg\min|\hat{D}_{c}^{\text{align}}-D_{c}|M_{c}$ (17)
In this way, the depth prediction can be continuously refined online with increasing observations, providing geometric supervisory signals during SLAM optimization. This helps guide global reconstruction and pose estimation, leading to improved results throughout the SLAM process.
III-D Semantic Gaussian Splatting SLAM
The global semantic 3D Gaussian Splatting SLAM system is illustrated in the left part of Fig. 2. Based on the established semantic 3D Gaussian representation, global 3D reconstruction (Mapping) and pose estimation (Tracking) are alternately performed with the video stream input. For the RGB-D setting, depth information for each image frame is obtained directly from depth sensors. For the monocular setting, we employ a feed-forward method to obtain the geometric prior, as introduced in the previous section. Further details of our Hier-SLAM++ system are provided in the following section.
III-D1 Semantic 3D Gaussian representation
The semantic 3D Gaussian representation is represented as the Gaussian primitives with the hierarchical semantic embedding. Motivated by [16], we adopt the isotropic and view-independent Gaussian integrated with the semantic embeddings. Each semantic Gaussian is composed of color $\bm{c}$, the center position $\bm{\mu}$, the radius $r$, the opacity $o$, and its semantic embedding $\bm{h}$. And the influence of each Gaussian $G(\bm{X})$ [13] according to the standard Gaussian equation is:
$G(\bm{X})=o\>\exp\left(-\frac{||\bm{X}-\bm{\mu}||^{2}}{2r^{2}}\right)$ (18)
where $\bm{X}$ stands for the 3D point.
Following [13], every semantic 3D Gaussian primitive within the current global map is projected to the 2D image space using the tile-based differentiable $\alpha$-compositing rendering. The semantic map $H$ is rendered as follows:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{H=\sum_{i=1}^{n}\bm{h}_{i}\alpha_{i}(\bm{X})T_{i}\quad\text{with}\quad T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j}(\bm{X}))}}$ (19)
The rasterized color image $C$, depth image $D$, and the silhouette image $S$ are defined as follows:
$C=\sum_{i=1}^{n}\bm{c}_{i}\alpha_{i}(\bm{X})T_{i},\;D=\sum_{i=1}^{n}\bm{d}_{i}\alpha_{i}(\bm{X})T_{i},\;S=\sum_{i=1}^{n}\alpha_{i}(\bm{X})T_{i}$ (20)
Different from previous work [16], which uses separate forward and backward Gaussian modules for different parameters, our Hier-SLAM++ system adopt a unified forward and backward modules that processes all parameters, including semantics, color, depth, and silhouette images—significantly improving overall runtime efficiency.
III-D2 Tracking
The tracking step focuses on pose estimation for each frame. For each incoming frame, the constant velocity model is first applied to initialize its pose based on the previous frame’s pose estimation. Following that, a pose optimization is performed with the global Gaussian map $G$ fixed:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathbf{T}_{c}=\arg\min_{\mathbf{T}_{c}}L_{\text{Track}}(G,\mathbf{T}_{c})}}$ (21)
where the tracking loss includes the rendering color and depth losses:
$L_{\text{{Track}}}=M\big{(}w_{1}L_{\text{{Depth}}}(G,\mathbf{T}_{c})+w_{2}L_{\text{{Color}}}(G,\mathbf{T}_{c})\big{)}$ (22)
where $L_{\text{Depth}}$ and $L_{\text{Color}}$ denote the L1-loss for the rendered depth and color information. The rendering follows the tile-based differentiable $\alpha$-compositing procedure, utilizing the current pose estimation along with the current global map estimation. Here, weights $w_{1}$ and $w_{2}$ are introduced to balance the two loss terms, and the optimization is only performed on the silhouette-visible image regions, defined as $M=(S>\delta)$. For the RGB-D setting, depth supervisory signals are directly obtained from depth sensors, whereas for the monocular setting, we leverage the geometric prior derived from the feed-forward method.
III-D3 Mapping
The global map information, including semantic information, is optimized in the mapping procedure while keeping camera poses fixed:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{G=\arg\min_{G}L_{\text{Map}}(G,\mathbf{T}_{i})}}$ (23)
The mapping optimization losses include depth, color, and semantic losses:
$L_{\text{{Map}}}=w_{3}ML_{\text{{Depth}}}(G,\mathbf{T}_{i})+w_{4}L_{\text{{Color}}}^{\prime}(G,\mathbf{T}_{i})+w_{5}L_{\text{{Semantic}}}(G,\mathbf{T}_{i})$ (24)
where $L_{\text{{Semantic}}}$ is the proposed semantic loss introduced in Section III-B, and $L_{\text{{Color}}}^{\prime}$ is the weighted sum of SSIM color loss and L1-Loss. The weights $w_{3}$, $w_{4}$, and $w_{5}$ are used to balance the different loss terms. Similar to the tracking process, depth supervisory signals come from sensors in the RGB-D setting. For the monocular setting, geometric priors derived from the feed-forward method are applied, and the two parameters (scale and shift) are further jointly optimized within the mapping process.
IV Experiments
IV-A Experiment settings
The experiments are conducted on both synthetic and real-world datasets, including ScanNet [40], Replica [41], and TUM-RGBD [42]. Following the evaluation metrics which used in previous SLAM methods [16, 33], we leverage ATE RMSE (cm) to quantify SLAM tracking accuracy. For mapping evaluation, Depth L1 (cm) is used to measure accuracy. To assess image rendering quality, we adopt PSNR (dB), SSIM, and LPIPS as evaluation metrics. In line with previous works [24, 7, 11], we assess semantic rendering performance using mIoU (mean Intersection over Union across all classes) to represent global semantic information. To evaluate computational efficiency, we report the runtime of the proposed SLAM method. For comparison, we benchmark our approach against state-of-the-art dense visual SLAM methods, including both NeRF-based and 3D Gaussian SLAM approaches, to demonstrate its effectiveness. Additionally, we compare with state-of-the-art semantic SLAM methods, covering both NeRF-based and Gaussian-based techniques, to highlight its capability in hierarchical semantic representation and scalability. The experiments are conducted in the Nvidia L40S GPU.
For experimental settings, the semantic embedding of each Gaussian primitive is initialized randomly. The semantic optimization loss weights $\omega_{1}$ and $\omega_{2}$ are set to $1.0$ and $0.0$, respectively, for the first $\eta$ iterations, where $\eta$ is set as $15$. Afterwards, $\omega_{1}$ remains $1.0$, while $\omega_{2}$ is increased to $5.0$. This strategy means that we first apply the Inter-level loss to initialize the hierarchical semantic coding, followed by incorporating the Cross-level loss to refine the whole embedding. For tracking loss, we use $\delta=0.99$, with weights $w_{1}=1.0$ and $w_{2}=0.5$. For mapping, the loss weights are set as follows: $w_{3}=1.0$, $w_{4}=0.5$, and $w_{5}=0.2$.
TABLE I: RGB-D Tracking performance ATE RMSE (cm) on the Replica.
Best results are highlighted as FIRST, SECOND.
Methods Avg. R0 R1 R2 Of0 Of1 Of2 Of3 Of4
iMap [32] 4.15 6.33 3.46 2.65 3.31 1.42 7.17 6.32 2.55
NICE-SLAM [33] 1.07 0.97 1.31 1.07 0.88 1.00 1.06 1.10 1.13
Vox-Fusion [43] 3.09 1.37 4.70 1.47 8.48 2.04 2.58 1.11 2.94
co-SLAM [44] 1.06 0.72 0.85 1.02 0.69 0.56 2.12 1.62 0.87
ESLAM [35] 0.63 0.71 0.70 0.52 0.57 0.55 0.58 0.72 0.63
Point-SLAM [45] 0.52 0.61 0.41 0.37 0.38 0.48 0.54 0.69 0.72
MonoGS-RGBD [17] 0.79 0.47 0.43 0.31 0.70 0.57 0.31 0.31 3.2
SplaTAM [16] 0.36 0.31 0.40 0.29 0.47 0.27 0.29 0.32 0.55
[…]
DNS SLAM [7] 0.45 0.49 0.46 0.38 0.34 0.35 0.39 0.62 0.60
SemGauss-SLAM [24] 0.33 0.26 0.42 0.27 0.34 0.17 0.32 0.36 0.49
SGS-SLAM [23] 0.41 0.46 0.45 0.29 0.46 0.23 0.45 0.42 0.55
Hier-SLAM [25] 0.33 0.21 0.49 0.24 0.29 0.16 0.31 0.37 0.53
[…]
Hier-SLAM++ (binary) 0.31 0.23 0.46 0.23 0.29 0.15 0.27 0.34 0.54
IV-B Camera Tracking Accuracy
Tab. I presents our tracking performance on the Replica dataset [41] in the RGB-D setting, compared with state-of-the-art dense SLAM methods, both with and without semantic information. By accurately learning semantic information, our Hier-SLAM++ outperforms all existing methods in 6 out of 8 sequences. For the remaining two sequences, our method achieves nearly the same performance with sub-millimeter level localization error. The lowest average tracking error clearly demonstrate that our approach surpasses other state-of-the-art methods. We present our RGB-D tracking performance on the real-world ScanNet dataset [40] in Tab. II. Compared to the results on the Replica dataset, all methods exhibit degraded performance due to the noisy and sparse depth sensor input, as well as lower color image quality caused by motion blur. Specifically for semantic understanding, the semantic annotations provided by ScanNet are significantly noisier than those in Replica, containing noticeable noise and imprecise boundaries. All of these issues negatively impact the overall performance of semantic SLAM systems. We evaluate all six sequences and show that our method achieves performance comparable to state-of-the-art approaches [25, 16], demonstrating the robustness of our proposed method under these challenging conditions. The tracking performance on the TUM-RGBD dataset is shown in Tab. III. Since the TUM-RGBD dataset does not contain semantic ground truth, we evaluate our method based on tracking accuracy here. Despite the low-quality RGB and depth inputs, as well as the absence of semantic information, all of which could negatively impact overall performance, our method still achieves similar performance to the state-of-the-art approach [16], with a localization error of less than $0.2$ cm, while operating at a significantly faster runtime. A detailed analysis of the runtime performance will be provided in the following section.
TABLE II: RGB-D Tracking performance ATE RMSE (cm) on the Scannet.
Best results are highlighted as first, second, third,fourth.
Methods Avg. $0$ 0059 0106 0169 0181 0207
NICE-SLAM [33] 10.70 12.00 14.00 7.90 10.90 13.40 6.20
Vox-Fusion [43] 26.90 68.84 24.18 8.41 27.28 23.30 9.41
Point-SLAM [45] 12.19 10.24 7.81 8.65 22.16 14.77 9.54
SplaTAM [16] 11.88 12.83 10.10 17.72 12.08 11.10 7.46
SGS-SLAM [23] 9.87 11.15 9.54 10.43 10.70 11.28 6.11
SemGauss-SLAM [24] 11.87 7.97 8.70 9.78 8.97
Hier-SLAM [25] 11.36 11.45 9.61 17.80 11.93 10.04 7.32
[…]
Hier-SLAM++ (binary) 11.59 13.24 9.60 17.83 11.75 9.83 7.29
TABLE III: RGB-D Tracking performance ATE RMSE (cm) on the TUM-RGBD. Best results are highlighted as first, second, third, fourth.
Methods Avg. fr1/desk fr1/desk2 fr1/room fr2/xyz fr3/off
Kintinuous [46] 4.84 3.70 7.10 7.50 2.90 3.00
ElasticFusion [47] 6.91 2.53 6.83 21.49 1.17 2.52
ORB-SLAM2 [29] 1.98 1.60 2.20 4.70 0.40 1.00
NICE-SLAM [33] 15.87 4.26 4.99 34.49 31.73 3.87
Vox-Fusion [43] 11.31 3.52 6.00 19.53 1.49 26.01
Point-SLAM [45] 8.92 4.34 4.54 30.92 1.31 3.48
SplaTAM [16] 5.48 3.35 6.54 11.13 1.24 5.16
[…]
IV-C Global Mapping Performance
Following previous methods [25, 23, 24], we evaluate the depth error on the Replica dataset [41] to assess overall mapping performance, as it provides the highest-quality depth maps for comparison. With an accurate understanding of semantics, our method effectively learns precise global geometry in unknown environments, benefiting both tracking and mapping in the semantic SLAM system. Our method achieves superior results in 7 out of 8 sequences and maintains a low average depth error across all sequences, with a difference of approximately $0.1$ cm compared to the best performing. These results demonstrate that our approach have strong semantic mapping capability, highlighting our superior mapping performance.
IV-D Rendering Quality
We present the quantitative analysis of the rendering performance of our proposed method in Tab. V. Following previous works, Point-SLAM [45] and NICE-SLAM [33], the evaluation is conducted on the input views. As shown in Tab. V, our method achieves superior image rendering performance compared to state-of-the-art approaches across all 8 sequences, as well as in the overall average evaluation. As discussed in previous sections, incorporating semantic information introduces a significant computational burden to the joint optimization process within the SLAM system. However, our method demonstrates notably improved rendering performance compared to all existing methods, including the non-semantic SLAM approaches. Although the additional computation for semantic learning increases the optimization complexity, the accurate semantic understanding enhances both geometric and appearance learning through joint optimization, effectively improving rendering performance.
The visual comparison of rendered images is presented in Fig. 6. Benefiting from precise semantic information learning and the improved SLAM system, the rendering results exhibit significantly improved detail preservation, such as more accurate and well-defined lamp boundaries in the first and fifth columns, and precise details even in fine structures like shutters in the second column. Moreover, our method produces high-quality renderings with fewer noise artifacts, as evidenced by the improved rendering quality of the floor in the fourth column and the chair in the sixth column. This demonstrates that better mapping performance leads to enhanced rendering quality. The joint optimization of semantic losses and photometric loss is mutually beneficial, facilitating more accurate global 3D reconstruction through semantic learning, which in turn improves the rendering performance of our Hier-SLAM++.
TABLE IV: Reconstruction performance Depth L1 (cm) on Replica. Best results are highlighted as first, second.
Methods Avg. R0 R1 R2 Of0 Of1 Of2 Of3 Of4
NICE-SLAM [33] 2.97 1.81 1.44 2.04 1.39 1.76 8.33 4.99 2.01
Vox-Fusion [43] 2.46 1.09 1.90 2.21 2.32 3.40 4.19 2.96 1.61
Co-SLAM [44] 1.51 1.05 0.85 2.37 1.24 1.48 1.86 1.66 1.54
ESLAM [35] 0.95 0.73 0.74 1.26 0.71 1.02 0.93 1.03 1.18
SNI-SLAM [11] 0.77 0.55 0.58 0.87 0.55 0.97 0.89 0.75 0.97
SGS-SLAM [23] 0.36
SemGauss-SLAM [24] 0.50 0.54 0.46 0.43 0.29 0.22 0.51 0.98 0.56
Hier-SLAM [25] 0.49 0.58 0.40 0.40 0.29 0.19 0.51 0.95 0.57
[…]
Hier-SLAM++ (binary) 0.49 0.59 0.40 0.40 0.30 0.17 0.52 0.95 0.57
IV-E Running time
The runtime results for the compared methods are provided in Tab. VI, covering both NeRF-based and Gaussian Splatting-based SLAM methods. Since the compared baseline using a flat representation (Hier-SLAM++-flat) cannot run successfully on an NVIDIA RTX 4090 due to GPU memory limitations, we report the runtime separately for RTX 4090 and L40S. In the first block of Tab. VI, we observe that, compared to NeRF-based methods [33], Gaussian-based state-of-the-art methods [16] achieve significantly higher operating speeds, benefiting from the fast rendering and efficient optimization capabilities inherent to the new 3D representation, i.e., Gaussian Splatting. Furthermore, with the carefully optimized forward and backward Gaussian rendering module introduced in our system, which serves as the core component of Gaussian Splatting-based SLAM systems, our method (Hier-SLAM++- w/o sem) achieves up to 2.4× faster tracking and 2.2× faster mapping compared to the best-performing method [16].
In the second block of Tab. VI, when integrating semantic information, our method maintains high efficiency by leveraging hierarchical semantic coding (Hier-SLAM++-one-hot), achieving nearly 3× faster tracking compared to flat semantic coding (Hier-SLAM++-flat). Compared to the state-of-the-art method [25], our Hier-SLAM++ attains similar operational speed but significantly outperforms in semantic segmentation accuracy. Thanks to the improved semantic decoder in our hierarchical semantic loss, our method achieves a $13\%$ mIoU improvement while incurring only an additional $0.04$ s per frame for tracking and $2$ s per frame for mapping. Further details on this improvement will be provided in the next section. Additionally, our binary version (Hier-SLAM++-binary) is able to achieve more efficient operation performance. Notably, our Hier-SLAM++ achieves a rendering speed of 2000 FPS, which increases to 3000 FPS when semantic information is not included.
TABLE V: Rendering performance PSNR, SSIM, LPIPS on Replica. Best results are highlighted as first, second.
Methods Metrics Avg. room0 room1 room2 office0 office1 office2 office3 office4
Visual SLAM
NICE-SLAM [33] PSNR $\uparrow$ 24.42 22.12 22.47 24.52 29.07 30.34 19.66 22.23 24.94
SSIM $\uparrow$ 0.809 0.689 0.757 0.814 0.874 0.886 0.797 0.801 0.856
LPIPS $\downarrow$ 0.233 0.330 0.271 0.208 0.229 0.181 0.235 0.209 0.198
Vox-Fusion [43] PSNR $\uparrow$ 24.41 22.39 22.36 23.92 27.79 29.83 20.33 23.47 25.21
SSIM $\uparrow$ 0.801 0.683 0.751 0.798 0.857 0.876 0.794 0.803 0.847
LPIPS $\downarrow$ 0.236 0.303 0.269 0.234 0.241 0.184 0.243 0.213 0.199
Co-SLAM [44] PSNR $\uparrow$ 30.24 27.27 28.45 29.06 34.14 34.87 28.43 28.76 30.91
SSIM $\uparrow$ 0.939 0.910 0.909 0.932 0.961 0.969 0.938 0.941 0.955
LPIPS $\downarrow$ 0.252 0.324 0.294 0.266 0.209 0.196 0.258 0.229 0.236
ESLAM [35] PSNR $\uparrow$ 29.08 25.32 27.77 29.08 33.71 30.20 28.09 28.77 29.71
SSIM $\uparrow$ 0.929 0.875 0.902 0.932 0.960 0.923 0.943 0.948 0.945
LPIPS $\downarrow$ 0.239 0.313 0.298 0.248 0.184 0.228 0.241 0.196 0.204
SplaTAM [16] PSNR $\uparrow$ 34.11 32.86 33.89 35.25 38.26 39.17 31.97 29.70 31.81
SSIM $\uparrow$ 0.968 0.978 0.969 0.979 0.977 0.978 0.969 0.949 0.949
LPIPS $\downarrow$ 0.102 0.072 0.103 0.081 0.092 0.093 0.102 0.121 0.152
Semantic SLAM
SNI-SLAM [11] PSNR $\uparrow$ 29.43 25.91 28.17 29.15 31.85 30.34 29.13 28.75 30.97
SSIM $\uparrow$ 0.921 0.884 0.900 0.921 0.935 0.925 0.930 0.932 0.936
LPIPS $\downarrow$ 0.237 0.307 0.292 0.265 0.185 0.211 0.230 0.209 0.198
SGS-SLAM [23] PSNR $\uparrow$ 34.66 32.50 34.25 35.10 38.54 39.20 32.90 32.05 32.75
SSIM $\uparrow$ 0.973 0.976 0.978 0.981 0.984 0.980 0.967 0.966 0.949
LPIPS $\downarrow$ 0.096 0.070 0.094 0.070 0.086 0.087 0.101 0.115 0.148
SemGauss-SLAM [24] PSNR $\uparrow$ 35.03 32.55 33.32 35.15 38.39 39.07 32.11 31.60 35.00
SSIM $\uparrow$ 0.982 0.979 0.970 0.987 0.989 0.972 0.978 0.972 0.978
LPIPS $\downarrow$ 0.062 0.055 0.054 0.045 0.048 0.046 0.069 0.078 0.093
Hier-SLAM [25] PSNR $\uparrow$ 35.70 32.83 34.68 36.33 39.75 40.93 33.29 32.48 35.33
SSIM $\uparrow$ 0.980 0.976 0.979 0.987 0.988 0.989 0.975 0.971 0.976
LPIPS $\downarrow$ 0.067 0.060 0.063 0.052 0.050 0.049 0.083 0.081 0.094
Hier-SLAM++ (one-hot) PSNR $\uparrow$ 35.61 32.76 34.64 36.27 39.65 40.12 33.46 32.49 35.52
SSIM $\uparrow$ 0.980 0.978 0.979 0.987 0.987 0.985 0.977 0.971 0.977
LPIPS $\downarrow$ 0.067 0.058 0.058 0.051 0.053 0.064 0.077 0.082 0.100
[…]
SSIM $\uparrow$ 0.981 0.975 0.979 0.988 0.987 0.990 0.978 0.971 0.976
LPIPS $\downarrow$ 0.066 0.064 0.061 0.049 0.056 0.043 0.076 0.081 0.098
TABLE VI: Runtime on Replica/R0. Best results are highlighted as first.
Methods Tracking Mapping Tracking Mapping
/Iteration /Iteration /Frame /Frame
RTX 4090 NICE-SLAM [33] 122.42 104.25 1.22 6.26
SplaTAM [16] 44.27 50.07 1.77 3.00
Hier-SLAM++ (w/o sem) 18.71 22.93 0.75 1.38
Hier-SLAM++ (one-hot) 37.63 194.78 1.50 11.69
L40S Hier-SLAM [25] 61.23 170.30 2.45 10.22
Hier-SLAM++ (flat) 168.94 204.25 6.75 12.26
Hier-SLAM++ (one-hot) 62.21 212.62 2.49 12.76
Hier-SLAM++ (binary) 58.91 210.65 2.36 12.64
  • Hier-SLAM++ (w/o sem) represents our proposed system without semantic information.
  • The units are as follows: Tracking/Iteration (ms), Mapping/Iteration (ms), Tracking/Frame (s), and Mapping/Frame (s).
TABLE VII: Semantic performance mIoU (%) and Parameter usage (MB) on Replica. Evaluations are conducted over a total of 102 semantic classes. Results are highlighted as first, second.
Methods Avg. R0 R1 R2 Of0
mIoU (%) NIDS-SLAM [12] 82.37 82.45 84.08 76.99 85.94
DNS-SLAM [7] 84.77 88.32 84.90 81.20 84.66
Hier-SLAM [25] 76.44 76.62 78.31 80.39 70.43
Hier-SLAM++ (flat) 90.35 91.21 90.62 89.11 90.45
Hier-SLAM++ (one-hot) 89.41 86.38 89.26 91.55 90.46
Hier-SLAM++ (binary) 81.27 82.77 73.87 89.64 78.80
SGS-SLAM [23] 92.72 92.95 92.91 92.10 92.90
SemGauss-SLAM [24] 96.34 96.30 95.82 96.51 96.72
SNI-SLAM [11] 87.41 88.42 87.43 86.16 87.63
Param (MB) Hier-SLAM [25] 910.50 793 1126 843 880
Hier-SLAM++ (flat) 2662.25 2355 3072 2560 2662
Hier-SLAM++ (one-hot) 927.50 814 1126 867 903
Hier-SLAM++ (binary) 591.75 528 690 563 586
  • The gray background indicates that the different evaluation methods are used, where evaluations are conducted on a subset of the total classes.
IV-F Hierarchical semantic understanding
We perform semantic understanding experiments on the synthetic dataset Replica [41] to demonstrate the comprehensive performance of our proposed method. Replica [41] is a synthetic indoor dataset that includes 102 semantic classes with high-quality semantic annotations. Similar to previous works [24, 7, 11], we evaluate the mIoU performance as the quantitative semantic results in Tab. VII, where mIoU (%) is reported by comparing the rendered semantic map with the ground truth across all original semantic classes (102 classes). To further analyze the storage requirements of semantic SLAM methods, we report the parameter storage usage in the second block of Tab. VII.
From Tab. VII, our method, along with the flat coding version and the proposed one-hot version, both demonstrate superior semantic performance compared to all existing methods. By utilizing hierarchical semantic representation, Hier-SLAM++ (one-hot) achieves a similar semantic performance with less than a $1\%$ difference in mIoU while requiring 4.5× less storage compared to the flat version. The training time shown in Tab. VI demonstrates that our hierarchical version (Hier-SLAM++-one-hot) requires only 37% of the tracking time per frame while maintaining similar frame mapping performance compared to the flat version (Hier-SLAM++-flat). The comparable mIoU performance, along with significantly reduced storage and computational cost, clearly demonstrates the effectiveness and efficiency of our proposed method. Furthermore, our binary version (Hier-SLAM++-binary) achieves state-of-the-art semantic understanding while achieves the best parameter storage efficiency, reducing storage usage to 63% of the one-hot version and requiring only 20% of the storage compared to the flat version. This reduction in storage usage highlights the efficiency of our proposed method. Because SGS-SLAM [23], SemGauss-SLAM [24] and SNI-SLAM [11] use different evaluation methods that restrict the mIoU calculation to a subset of the global semantic classes. Specifically, SGS-SLAM computes the average IoU for the semantic classes visible in each frame, and SemGauss-SLAM together with SNI-SLAM [11] evaluates performance based on 52 visible classes across all evaluation sequences, making direct comparisons unfair.
[…]
Overall, our method demonstrates superior performance in semantic understanding while significantly reducing storage requirements and training time, benefiting from the proposed hierarchical semantic representation.
Figure 4: Visualization of our semantic rendering performance on the Replica [41] dataset. The first four rows demonstrate rendered semantic segmentation in a coarse-to-fine manner. The fifth row exhibits the finest semantic rendering, equivalent to the flat representation with $102$ original semantic classes from the Replica dataset. The last row visualizes the semantic ground truth for comparison.
IV-G Scaling up capability
To showcase the scaling-up capability of Hier-SLAM++, we evaluate it on the real-world complex dataset, ScanNet [40], which contains up to 550 distinct semantic classes. In contrast to Replica [41], where the semantic ground truth is synthesized from a global world model and can be considered ideal, the semantic annotations in ScanNet are significantly noisier, containing considerable noise and incorrect or unclear class boundaries. Moreover, the dataset contains noisy depth sensor inputs and blurred color images, making semantic understanding particularly challenging in this scenes.
Using the flat semantic representation cannot even run successfully due to memory limitations. In contrast, we establish the hierarchical tree with the assistance of LLMs and 3D Generative Models, as introduced in Section-III, which guides the compaction of the coding from the original 550 semantic classes to 73 semantic symbolic codings, resulting in over 7 times reduction in coding usage. As visualized in Fig. 5, our estimated 3D global semantic map at different levels demonstrates a coarse-to-fine semantic understanding, showcasing our method’s scaling-up capability in handling real-world complex scenes.
TABLE VIII: Monocular quantitative results on the TUM RGB-D dataset. Best results are highlighted as first, second.
Sequences fr1-desk fr2-xyz
Loop Method RMSE (cm) $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ RMSE (cm) $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
w/ DROID-SLAM [48] 1.80 - - - 0.50 - - -
ORB-SLAM2 [29] 2.00 - - - 0.60 - - -
Photo-SLAM [19] 1.54 20.97 0.743 0.228 0.98 21.07 0.726 0.166
[…]
MonoGS [17] 4.15 21.02 0.703 0.362 4.79 22.21 0.728 0.288
Hier-SLAM++ 6.20 21.91 0.836 0.269 3.22 24.67 0.946 0.110
IV-H Monocular performance
We present our monocular SLAM performance in Tab. VIII. For camera tracking, our method achieves competitive performance compared to state-of-the-art approaches without loop closure. Specifically, compared to the state-of-the-art Gaussian Splatting-based methods [17], our method exhibits similar tracking performance while achieving superior rendering quality. For SLAM methods with loop closure, which typically demonstrate superior camera tracking performance due to well-executed loop closure mechanisms that significantly enhance tracking accuracy and overall SLAM performance, our method still achieves superior rendering accuracy compared to Photo-SLAM [19], which incorporates loop closure. This highlights the strong rendering capabilities of our proposed method. With the assistance of the geometric prior provided by the feed-forward method, the joint optimization of both photometric loss and aligned depth loss contributes to improved rendering performance in the whole SLAM system.
Figure 5: Visualization of the established semantic 3D map across multiple levels, illustrating a coarse-to-fine semantic understanding of the complex scene. The bottom of the figure presents localization, rendering, and depth performance, offering a comprehensive overview of Hier-SLAM++’s effectiveness and demonstrating the scaling-up capability of our proposed method.
IV-I Ablation study
We evaluate SLAM performance under different tree designs in Tab. IX. The evaluation is conducted using the baseline method, Hier-SLAM [25], ensuring that any performance differences are solely attributable to the tree structures. In Tab. IX, each metric represents the average result from two runs on the first four sequences—Room0, Room1, Room2, and Office0—from the Replica dataset. We generate two different 5-level trees using our proposed tree generation procedure for the Replica dataset and incorporate them into the semantic SLAM system. As shown in Tab. IX, all methods demonstrate good performance, which can be attributed to the well-designed tree structures generated through both LLMs and 3D Generative Models. Additionally, the fully optimized hierarchical semantic loss contributes to the robustness of SLAM performance. Among all compared methods, tree-1 achieves the best semantic mIoU performance, the highest depth mapping accuracy, and close rendering quality (PSNR, SSIM, and LPIPS) compared with other settings. Therefore, we select tree-1 as the hierarchical tree structure leveraged within our semantic SLAM system.
TABLE IX: Performance of different trees on Replica. Best results are highlighted in bold.
[…]
Hier-SLAM [25] 75.517 0.292 0.416 35.885 0.982 0.057
[…]
Hier-SLAM++${}^{\text{tree2}}$ 75.061 0.280 0.413 35.792 0.982 0.057
V Conclusion
We introduced Hier-SLAM++, a neuro-symbolic semantic 3D Gaussian Splatting SLAM system that supports both RGB-D and monocular inputs, incorporating a novel hierarchical categorical representation. Specifically, we proposed a compact and generalized hierarchical representation that encodes both semantic meaning and geometric attributes (i.e., size and shape), leveraging the capabilities of LLMs and 3D generative models. To enhance global semantic understanding, we introduced an improved hierarchical semantic loss. Furthermore, we extended the system from an RGB-D-only SLAM to support monocular input, utilizing geometric priors derived from a feed-forward SLAM approach. Experimental results demonstrate that Hier-SLAM++ achieves superior or on-par performance with state-of-the-art NeRF-based and Gaussian-based SLAM methods in terms of tracking, mapping, and semantic understanding accuracy, while significantly reducing storage requirements and runtime. These advances establish Hier-SLAM++ as a robust solution for semantic 3D understanding across diverse environments.
Figure 6: Visualization of RGB image rendering performances on the Replica dataset [41]. The first row presents the image rendering results of our Hier-SLAM++. The second to fourth rows show the rendering outputs of the comparison methods. The last row displays the ground truth images for reference. We use red boxes to highlight key differences regions within the image. The visual comparison clearly demonstrates that our method achieves superior rendering quality, closely resembling the ground truth images compared to state-of-the-art approaches.
APPENDIX
I Appendix 1. Implementation Details for Tree Generation
During the hierarchical tree generation process, we utilize LLMs to group flat semantic classes at different levels based on physical size, semantic function, and to summarize and refine geometric grouping, ensuring balanced, meaningful, and descriptive outputs. After generating each tree level, we further employ LLMs as validators to examine the results at each level, ensuring that all input categories are fully included with no omissions and that no extra categories are mistakenly introduced.
Size Grouping: The prompt used for the initial grouping based on object physical sizes is illustrated as follows:
Code I.1: Prompt for Size Grouping You are a smart assistant tasked with dividing the following items into meaningful groups based on their properties. The key requirements are:
Balance: The number of items in each group should be as evenly distributed as possible. A difference of 1 item between groups is acceptable, but larger differences should be avoided.
Meaningfulness: The groups must be meaningful, based on the items’ inherent characteristics. Grouping should make logical sense to a human observer.
Descriptive Group Names: Each group must have a clear and descriptive name that reflects its characteristics.
Goal: Cluster items into size groups, based on their typical physical size in scenes. The groups must strictly be by sizes, for example: small, small medium, medium, large, extra large, etc.
Items: {classes input}
Ensure groups are meaningful and provide descriptive group names. Output must follow this JSON format:
    "<GROUP_1>": ["<ITEM_1>", ...],
    "<GROUP_2>": ["<ITEM_2>", ...],
Function Grouping: The prompt adopted for the subsequent semantic function grouping is as follows:
Code I.2: Prompt for Function Grouping You are a smart assistant tasked with dividing the following items into meaningful groups based on their properties. The key requirements are:
Balance: The number of items in each group should be as evenly distributed as possible. A difference of 1 item between groups is acceptable, but larger differences should be avoided.
Meaningfulness: The groups must be meaningful, based on the items’ inherent characteristics. Grouping should make logical sense to a human observer.
Descriptive Group Names: Each group must have a clear and descriptive name that reflects its characteristics.
Goal: Cluster items in the ’size’ size group into functionality-based groups. The name of the clusters should not be too specific, it could be as general like storage, furniture, etc. Ensure that items in each group serve similar purposes or have similar functionalities.
Items: {classes input}
Ensure groups are meaningful and provide descriptive group names. Output must follow this JSON format:
    "<GROUP_1>": ["<ITEM_1>", ...],
    "<GROUP_2>": ["<ITEM_2>", ...],
Summarize and Refine for Geometric Grouping: After grouping with the 3D Generative Model, we employ LLMs to summarize the clustering results with descriptive labels and balance the nodes within each group. The prompt for LLMs is illustrated as follows:
Code I.3: Prompt for Geometric Grouping Results The following groups of indoor scene items have been clustered based on their shape: {formatted_clusters}
Instructions: Balance: Ensure the groups are as evenly distributed as possible. A difference of 1 item between groups is acceptable. If needed, move a small number of items from one group to another to achieve balance, feel free to even remove a group if it has only one member (just dont leave any groups empty), ensuring that the groups remain meaningful.
Meaningful Naming: After balancing the groups, assign a descriptive and meaningful name to each group, based on the shared shape characteristic. The name should clearly reflect the shape or geometric property of the items in that group. Always prefer singular shapes names like box, rectangle, soft, flat, etc.
No Dublications: make sure to not repeat any class members, the group names are fine but not the groups’ members.
The group names must strictly by shapes and DO NOT leave any group empty, you could remove it if its empty.
The output must be the same JSON format as below:
    "<GROUP_1>": ["<ITEM_1>", ...],
    "<GROUP_2>": ["<ITEM_2>", ...],
LLM Validators: The prompt used for LLM validator is as following:
Code I.4: Prompt for LLM validator The following groups of items in a scene have been clustered based on their shape: {formatted_clusters}
Instructions: Balance: Ensure the groups are as evenly distributed as possible. A difference of 1 item between groups is acceptable. If needed, move a small number of items from one group to another to achieve balance, remove a group if it has only one member, ensuring that the groups remain meaningful.
Meaningful Naming: After balancing the groups, assign a descriptive and meaningful name to each group, based on the shared shape characteristic. The name should clearly reflect the shape or geometric property of the items in that group. Always prefer singular shapes names like box, rectangle, soft, flat, etc.
New Group: If it is necessary to create a new group, feel free to do so, DO NOT create any non-original items.
No Dublications: make sure to not repeat any class members, the group names are fine but not the groups’ members.
The output must be the same JSON format as below:
    "<GROUP_1>": ["<ITEM_1>", ...],
    "<GROUP_2>": ["<ITEM_2>", ...],